Grade 12General Physics

A resistor ‘R’ and 2µF capacitor in series is connected through a switch to 200 V direct supply. Across the capacitor is a neon bulb that lights up at 120 V. Calculate the value of R to make the bulb light up 5 s after the switch has been closed

sudhanshu
12 Years agoGrade 12

1 Answer

Jayant gehlot
7 Years ago
q=cv(1-e-t/RC) charging of capacitor formula
q/c=V(1-e-t/RC)
v=V(1-e-t/RC)
120=200(1-e-t/RC)
e-t/RC=80/100
-t/RC=log(1/2.5)
t/RC=log(2.5)
R=t/C*log2.5
R=5/2*10-6*0.4
R=2.7*106